1. Suppose T: R3 -> R3 is linear and has an upper-triangular matrix with respect to the basis (1, 0, 0), (1,2,1), (1,2,2). Then the orthonormal basis of R3 with respect to which T has an upper-triangular matrix is...
  2. For u = (1,2,2) and v= (1, -2, -1), the value of ||u-v|| is...
  3. Suppose that u,v E V, where V is a real vector space such that ||u|| = 4 and ||v|| = 3 Then < u + v, u - v > is...
  4. In R3, let U = Span ((1, 0, 0), (0, 1/sqrt2 , 1/sqrt2 )) The u E U such that || u - ( 2,4,6)|| is as small as possible is...
  5. Let T: R3 -> R3 defined as T(x,y,z) = (2x, x+y, x-z). Then adjoint operator T* (u,v,w) is...

(i) (2u+v+w, u, -w)

(ii) (2u,v+w, u-w)

(iii) (u,v, -w)


1
Expert's answer
2021-10-28T15:30:01-0400

Section 1


Using Gram- Schmidt principle

"e_1=(1,0,0)"

"f_2=(1,2,1)-\frac{(1,2,1).(1,0,0)}{(1,0,0).(1,0,0)}(1,0,0)"

="(0,2,1)"

"e_2=\frac{f_2}{||f_2||}=\frac{(0,2,1)}{√5}"


"e_2=(\frac{0}{√5},\frac{2}{√5},\frac{1}{√5})"


"f_3=(1,2,2)-[(1,2,2).(1,2,2)](1,0,0)"

"-[(1,2,2).(0,2,1)](0,2,1)"

"=(0,-\frac{2}{5},\frac{4}{5})"


"\frac{(0,\frac{-2}{5},\frac{4}{5})}{||f_3||}"


="\frac{(0,\frac{-2}{5},\frac{4}{5})}{\frac{2√5}{5}}"


e3 ="(0,\frac{-1}{√5},\frac{2}{√5})"



Part 2

u-v=(1-1, 2- -2, 2- -1)=(0,4,3)

||u-v||= "\sqrt{\smash[b]{4^2+3^2}}=5"



Part 3

<u+v, u-v> = ||u||2 - 2 ||u|| ||v|| +||v||2

= 42 - 2(4)(3) +32

= 1


Part 4

let "u" = "x\begin{pmatrix}
 1 \\
 0 \\
0
\end{pmatrix}" "+ y\begin{pmatrix}
 0 \\
 \frac{1}{√2} \\
\frac {1}{√2}
\end{pmatrix}"



"=\begin{pmatrix}
 x \\
 \frac {y}{√2}\\
\frac {y}{√2}
\end{pmatrix}"


"||(x,\frac {y}{√2},\frac {y}{√2}) -(2,4,6)||"

"is \> minimum"


"(x-2)^2+(\frac{y}{√2}-4)^2+(\frac {y}{√2}-6)^2" "is \> minimum"


"x^2-4x+y^2-\frac{20y}{√2}+


\>56" "\>is \> minimum"


"\frac{d}{dx}(x^2-4x)=0"

"\implies 2x-4=0\quad x=2"


"\frac {d}{dy}(y^2-\frac{20y}{√2})=0"

"2y-\frac {20}{√2}=0\implies\>y=\frac{10}{√2}"


Substituting the value of "x \>and \>y \>in \>u"


"u=\begin{bmatrix}
 2 \\
 5 \\
5
\end{bmatrix}"


Part 5

Transforming matrix "T" ;

"\begin{pmatrix}
 2x \\
 x+y \\
x-z
\end{pmatrix}" ="x\begin{pmatrix}
 2\\
1\\ 
1
\end{pmatrix}" "+y\begin{pmatrix}
 0 \\
 1\\ 
0
\end{pmatrix}" +"z"


"\begin{pmatrix}
 0 \\
 0\\
-1
\end{pmatrix}"



"T=\begin{pmatrix}
 2& 0&0\\
 1&1 & 0\\
1&0&-1
\end{pmatrix}"



"T^*=T^T=\begin{pmatrix}
 2&1 & 1 \\
 0&1& 0\\
0&0&-1
\end{pmatrix}"


"\begin{pmatrix}
 2&1& 1\\
 0&1 & 0\\
0&0&-1
\end{pmatrix}\begin{pmatrix}
 u \\
 v\\
w
\end{pmatrix}" =


"\begin{pmatrix}
 2u+v+w \\
 v \\
-w
\end{pmatrix}"



Therefore the adjoint operator "T^*(u,v,w)"

"=(2u+v+w,u,-w)"